OpenStudy (anonymous):

Aki's Bicycle Designs has determined that when x hundred bicycles are built, the average cost per bicycle is given by C(x)=0.2x^2-1.6x+4.752, where C(x) is in hundreds of dollars. How many bicycles should the shop build to minimize the average cost per bicycle?

OpenStudy (anonymous):

I keep getting 112

OpenStudy (anonymous):

answer was 400

OpenStudy (anonymous):

Differentiate and set equal to zero for minimum.

OpenStudy (anonymous):

Like this 0.4x-1.6 = 0 -> x= 4

OpenStudy (anonymous):

did you try finding the vertex of the parabola 0.2x^2-0.6x+0.711 ?

OpenStudy (anonymous):

for y=ax^2+bx+c, the vertex is at x=-b/(2a)

OpenStudy (anonymous):

can you continue from here ?